OpenStudy (anonymous):

Can anyone help me with this? A model of seasonal growth is given by dx/dt=rx cos(wt) where r and w are constants. Illustrate the behaviour of the solution x of this equation.

OpenStudy (anonymous):

\[\frac{ dx }{ dt } =r*x *\cos(wt)\] as r and w are constants then we can apply product rule on the eq \[\frac{ dx }{ dt } = (r*x) *(\cos(wt))\] (r*x) = f(x) cos(wt) = g(x) \[\frac{ dx }{ dt } = f(x)*g \prime(x) + f \prime(x) *g(x)\]

OpenStudy (anonymous):

This is what i got from the equation.\[x=x_0 e ^{\frac{ r }{ w } \sin(wt)}\]But now, how should I explain about the behaviour of this solution?